# General diffraction properties of aperiodic slit arrays

### v1

@@ 1 Introduction
In contrast, diffraction by aperiodic gratings remains less explored. Prior studies have examined their ability to deflect light [6], their potential for measuring the radii of Gaussian beams [7], the focusing properties of chirped gratings [8], the potential advantages of Fibonacci gratings in image-forming devices [9], as well as the self-imaging capacities of fractal [10, 11] and quasiperiodic [12] gratings. X-ray diffraction by aperiodic solids such as quasicrystals, which present long-range order without translational symmetry, has also been a subject of considerable interest in crystallography [13, 14]. For instance, previous studies have investigated wave dynamics in photonic quasicrystals [15], the optical properties of deterministic aperiodic nanostructures [16], and light transmission through self-similar aperiodic structures [17].
@@ 3.1 Conditions for experimental observation
It is now useful to analyse the implications of this condition for the grating’s geometry. For this purpose, let us suppose that we want to design the diffraction grating such that Eq. (15) is exactly satisfied for $j=1$. To better understand this condition, let us first notice that for the case of a periodic grating, the condition implies the use of $s=L$, transforming the slit array into a completely opened aperture. We therefore conclude that, as long as the grating remains a periodic array of separate slits, at least two diffraction peaks would be visible at the transverse positions $x^{\prime}=\pm L^{\prime}_{1}$.
@@ 3.2 Experimental observation of the 0-th order periodicity
Let $\frac{1}{2}\Psi_{r}(x)$ denote the near-field structure on the right-hand side of $x=0$, so that $\Psi(x)=\frac{1}{2}[\Psi_{r}(x)+\Psi_{r}(-x)]$. From Eq. (11) it is easy to see that $\tilde{\Psi}_{r}(-x^{\prime})=\tilde{\Psi}_{r}^{*}(x^{\prime})$, where (*) stands for the complex conjugation operation. As a result, the far-field can be expressed as $\tilde{\Psi}(x^{\prime})=\frac{1}{2}\left[\tilde{\Psi}_{r}(x^{\prime})+\tilde{\Psi}_{r}^{*}(x^{\prime})\right]=\Re{[\tilde{\Psi}_{r}(x^{\prime})]}$, with $\Re$ denoting the real part of the function:
@@
with $L_{0}^{\prime}$ now appearing in the sum inside the argument of the cosine, instead of as a global phase.
In the next section, we shall use Eq. (17) to represent the Fraunhofer diffraction pattern of mirror-symmetric fields experimentally generated through aperiodic slit arrays.
@@ 4 Experimental results
Figure 1: Fig. (a) illustrates a simplified version of our experimental setup. A laser beam ($\lambda=660$nm) is sent to a Spatial Light Modulator. The unmodulated light is filtered out, and a CMOS camera records the Fraunhofer diffraction pattern of the modulated light. We also present images of the gratings for a slit distribution described by Eq. (18), with $L=(0.560\pm 0.008)$mm and slit widths of (b) s=0.024s=0.024mm (3 pixels of the SLM); (c) s=0.040s=0.040mm (5 pixels of the SLM); (d) s=0.056s=0.056mm (7 pixels of the SLM). The Gaussian illumination in the outlined regions is shown on the right side of the figure for $\bar{n}=25$, and $\sigma=2\sqrt{2}$. The position of the centre of each slit is represented by a vertical line.
@@
We have experimentally investigated the Fraunhofer diffraction pattern of aperiodic slit arrays. In our setup, a fibre-coupled laser beam with $\lambda=660$nm [Thorlabs S1FC660] is first expanded and sent to a spatial light modulator (SLM). The SLM [HoloEye Pluto NIR-015] is programmed to generate the field given in Eq. (1) that is then Fourier-transformed using a lens system with effective focal length f=200f=200mm. The Fraunhofer diffraction pattern is then registered using a CMOS imaging sensor [Basler ace U acA5472-17um]. A sketch of our experimental setup is shown in Fig. 1-(a).
The chosen slit array has an aperiodic slit distribution obeying
$x_{\pm n}=\pm L\sqrt{n},$ (18)
and the amplitude distribution of the light going through the slit array is chosen to be the Gaussian function given in Eq. (8) centred around two opposite values of the slit number: $\bar{n}$ and $\bar{n}$. Figs. 1-(b)-(d) present examples of gratings with slit positions following Eq. (18) (on the left), alongside their respective Gaussian intensity distributions (on the right). At first glance, the slit distributions of the illuminated parts may appear periodic, as the aperiodicity becomes more noticeable when analysing a larger portion of the gratings. However, the following experimental results show that the far-field diffraction patterns differ significantly from those of periodic gratings.
@@
Table 1: Far-field periodicities produced by the slit array given in Eq. (18) and experimental results obtained from the fitting of the measurements shown in Fig. 2
$L^{\prime}_{0}$ $L^{\prime}_{1}$ $L^{\prime}_{2}$
Symbolic equations for $L_{j}^{\prime}$ using the slit distribution of Eq. (18)
Eq. (18) $({2\pi}/{\xi L})\bar{n}^{-0.5}$ $({2\pi}/{\xi L})2\bar{n}^{0.5}$ $({2\pi}/{\xi L})8\bar{n}^{1.5}$
Values of $L_{j}^{\prime}$ calculated from the experimental parameters (mm)
Eq. (18) $0.0369\pm 0.0007$ $48\pm 0.01 11$$18\pm 1$
Values obtained from fitting of the experimental results (mm)
Fig. 2-(a) $1.49\pm 0.01$
Fig. 2-(b) $1.48\pm 0.01$
Fig. 2-(c) $0.0370\pm 0.0007$ $1.47\pm 0.01$
@@
Device resolution, such as the pixel size of modulators and cameras, sets the utmost limit when structuring or measuring light fields. For example, in our experimental setup, the distance between any pair of slits is resolution-limited by the pixel size of our SLM (we use the pixel size as the uncertainty for the length parameter $L$ of the slit array produced). Unlike this experimental limitation, in Subsection 3.1 we have shown that designing an arbitrary aperiodic array with geometry satisfying Eq. (15) for $j=1$ leads to a fundamental resolution limitation. The suppression of the maxima associated with the first-order periodicity by the zeros of the sinc function requires the width of the slits to be greater than the distance between some of the illuminated slits in the array. This condition implies that the part of the array with merged slits loses the aperiodic structure that is responsible for the Fraunhofer diffraction patterns studied in this paper. The resulting aperture is then described by a set of separate slits distributed aperiodically and a large aperture produced by a number of merged slits.
@@
Figure 3: Experimental results obtained using a grating with slit distribution given by Eq. (18) with $(L=0.400\pm 0.008)$mm and slit width s=0.04s=0.04mm (corresponding to 5 pixels of our SLM). The images correspond to the Fraunhofer diffraction patterns obtained by adjusting the illumination coefficients $A_{n}$ as the Gaussian function, Eq. (8), with $\sigma=\sqrt{2}$ and (a) $\bar{n}=\pm 10$, for (b) $\bar{n}=\pm 17$, and (c) $\bar{n}=\pm 25$. In (d), (e) and (f), we plot the graphs resulting from the sum of recorded images (a), (b) and (c), respectively, over 80 pixels in the vertical direction. Each graph was normalised so that the area under the curve equals one and then divided by the maximum value among all three graphs. The black-dashed line corresponds to the plot of the sinc function described by Eq. (18).
Another way to approach the upper limit of (19) is by increasing the width of the slit $s$ while maintaining fixed values for both $L$ and $\bar{n}$. In this case, $L_{1}^{\prime}$ remains constant, while the positions of the zeros of the sinc function decrease. In Fig. 4 we present the experimental results for a grating with $L=(0.560\pm 0.008)$$m$, $\bar{n}=\pm 25$ and s=0.024s=0.024mm [Fig. 4-(a)], s=0.040s=0.040mm [Fig. 4-(b)] and s=0.056s=0.056mm [Fig. 4-(c)], with the third case corresponding to total suppression of the diffraction maxima at $x^{\prime}=\pm L_{1}^{\prime}$. The respective sums over the vertical direction are shown in Figs. 4-(d), -(e) and -(f). In this case, the reduction of the lateral maxima causes the central peak to become visibly taller.
Figure 4: Experimental results obtained using a grating with slit distribution given by Eq. (18) with $L=(0.560\pm 0.008)$$m$, a Gaussian illumination distribution with $\bar{n}=\pm 25$ and $\sigma=2\sqrt{2}$ The images correspond to the Fraunhofer diffraction patterns obtained with three different slits widths: (a) s=0.024s=0.024mm (3 pixels of the SLM); (b) s=0.040s=0.040mm (5 pixels of the SLM); (c) s=0.056s=0.056mm (7 pixels of the SLM). In (d), (e) and (f), we plot the graphs resulting from the sum of recorded images (a), (b) and (c), respectively, over 80 pixels in the vertical direction. Each graph was normalised so that the area under the curve equals one and then divided by the maximum value among all three graphs. The dashed black line corresponds to the plot of the sinc function described by Eq. (18).
@@ 5 Conclusions
We have theoretically and experimentally investigated the Fraunhofer diffraction patterns produced by aperiodic slit arrays. We provide general conditions for the observation of quasiperiodic oscillations over different length scales based on the geometry and illumination pattern of the array. These results provide new insights into the diffraction properties of aperiodic gratings, with possible applications in beam shaping, the design of diffractive devices, and the characterisation of materials with quasiperiodic structures.
@@
Data underlying the results presented in this paper are not publicly available but may be obtained from the authors upon reasonable request.

### v2

@@ 1 Introduction
In contrast, diffraction by aperiodic gratings remains less explored. Prior studies have examined their ability to deflect light [6], their potential for measuring the radii of Gaussian beams [7], the focusing properties of chirped gratings [8], the potential advantages of Fibonacci gratings in image-forming devices [9], as well as the self-imaging capacities of fractal [10, 11] and quasiperiodic [12] gratings. Aperiodic gratings have also been employed in the generation of achromatic [13] and single-photon [14] super-oscillations. X-ray diffraction by aperiodic solids such as quasicrystals, which present long-range order without translational symmetry, has also been a subject of considerable interest in crystallography [15, 16]. For instance, previous studies have investigated wave dynamics in photonic quasicrystals [17], the optical properties of deterministic aperiodic nanostructures [18], and light transmission through self-similar aperiodic structures [19].
@@ 3 Far-field diffraction periodicities of general slit arrays
The intensity of the Fraunhofer diffraction pattern is obtained by taking the squared modulus of the amplitude given in Eq. (11):
$I(x^{\prime})=|\tilde{\Psi}(x^{\prime})|^{2}=\tilde{\Pi}^{2}(x^{\prime})\sum_{n}\sum_{m}A_{n}A_{m}\exp\left[-i\sum_{j=1}^{\infty}\frac{2\pi}{L^{\prime}_{j}}(n-\bar{n})^{j}x^{\prime}\right]\exp\left[i\sum_{j=1}^{\infty}\frac{2\pi}{L^{\prime}_{j}}(m-\bar{n})^{j}x^{\prime}\right]\;,$ (15)
which can be reduced, using the symmetry under the exchange $n\leftrightarrow m$, to:
$I(x^{\prime})=\tilde{\Pi}^{2}(x^{\prime})\sum_{n}\sum_{m}A_{n}A_{m}\cos{\Bigg\{\sum_{j=1}^{\infty}\frac{2\pi}{L^{\prime}_{j}}\left[(n-\bar{n})^{j}-(m-\bar{n})^{j}\right]x^{\prime}\Bigg\}}\;.$ (16)
As in the case of the amplitude distribution, Eq. (11), the parameters $L_{j}^{\prime}$ with $j\geq 1$ also appear as periodicities in the intensity pattern, Eq. (16). By contrast, the 0-th order periodicity ($L_{0}^{\prime}$) is not directly observable in intensity. The experimental observation of this lowest order periodicity is discussed in Subsection 3.2.
@@ 3.1 Conditions for experimental observation
It is now useful to analyse the implications of this condition for the grating’s geometry. For this purpose, let us suppose that we want to design the diffraction grating such that Eq. (17) is exactly satisfied for $j=1$. To better understand this condition, let us first notice that for the case of a periodic grating, the condition implies the use of $s=L$, transforming the slit array into a completely opened aperture. We therefore conclude that, as long as the grating remains a periodic array of separate slits, at least two diffraction peaks would be visible at the transverse positions $x^{\prime}=\pm L^{\prime}_{1}$ within the zeroth diffraction order imposed by the slit width.
@@ 3.2 Experimental observation of the 0-th order periodicity
Let $\frac{1}{2}\Psi_{r}(x)$ denote the near-field structure on the right-hand side of $x=0$, so that $\Psi(x)=\frac{1}{2}[\Psi_{r}(x)+\Psi_{r}(-x)]$. From Eq. (11) it is easy to see that $\tilde{\Psi}_{r}(-x^{\prime})=\tilde{\Psi}_{r}^{*}(x^{\prime})$, where the asterisk denotes the complex conjugation operation. As a result, the far-field can be expressed as $\tilde{\Psi}(x^{\prime})=\frac{1}{2}\left[\tilde{\Psi}_{r}(x^{\prime})+\tilde{\Psi}_{r}^{*}(x^{\prime})\right]=\Re{[\tilde{\Psi}_{r}(x^{\prime})]}$, with $\Re$ denoting the real part of the function:
@@
with $L_{0}^{\prime}$ now appearing in the sum inside the argument of cosine, instead of as a global phase.
In the next section, we shall use Eq. (19) to represent the Fraunhofer diffraction pattern of mirror-symmetric fields experimentally generated through aperiodic slit arrays. In order to observe the diffraction features discussed in this work, we consider a slit distribution for which the different-order periodicities differ in scale, that is, satisfying the condition in Eq. (13). Furthermore, the distance between neighbouring slits in the chosen distribution decreases with the slit index, allowing for the suppression of first-order maxima through an appropriate choice of parameters. In particular, the chosen array obeys the following slit distribution
$x_{\pm n}=\pm L\sqrt{n}\;.$ (20)
The periodicities of the far-field diffraction pattern can be determined by substituting the distribution given in Eq. (20) into Eq. (12). The corresponding zeroth-, first-, and second-order periodicities are given by
$L^{\prime}_{0}$ $=\frac{2\pi}{\xi L}\bar{n}^{-\frac{1}{2}}\;,$ (21)
$L^{\prime}_{1}$ $=\frac{2\pi}{\xi L}2\bar{n}^{\frac{1}{2}}\;,$ (22)
$L^{\prime}_{2}$ $=\frac{2\pi}{\xi L}8\bar{n}^{\frac{3}{2}}\;.$ (23)
Let us examine the conditions under which these periodicities are well separated in scale. The ratios between the successive order periodicities are $\frac{L^{\prime}_{1}}{L^{\prime}_{0}}=2\bar{n}$ and $\frac{L^{\prime}_{2}}{L^{\prime}_{1}}=4\bar{n}$. Thus, the condition in Eq. (13) is satisfied when $\bar{n}\gg 1$.
Finally, it is worth remarking that, for a sufficiently smoothly varying amplitude $A_{n}$, it is possible to obtain approximate analytical descriptions of the sum in Eq. (19), each of which is valid for a separate diffraction peak, by using the Poisson summation formula [21] and making a stationary-phase approximation. See, e.g., Refs. [22], [23] for details, and Ref. [20] for an application to the diffraction structure described by Eq. (20).
@@ 4 Experimental results
Figure 1: (a) Simplified version of our experimental setup. A laser beam ($\lambda=660$nm) is sent to a Spatial Light Modulator (SLM). The unmodulated light is filtered out, and a Complementary Metal-Oxide-Semiconductor (CMOS) camera records the Fraunhofer diffraction pattern of the modulated light. We also present images of the aperiodic gratings for a slit distribution described by Eq. (20), with $L=(0.560\pm 0.008)$mm and slit widths of (b) s=0.024s=0.024mm (3 pixels of the SLM); (c) s=0.040s=0.040mm (5 pixels of the SLM); (d) s=0.056s=0.056mm (7 pixels of the SLM). The Gaussian illumination in the outlined regions is shown on the right-hand side of the figure for $\bar{n}=25$, and $\sigma=2\sqrt{2}$, with the position of the centre of each slit represented by a vertical line. Note that the region of the array where the slits are merged increases as the slits get wider. The parameters used to draw parts (b), (c) and (d) of this figure correspond to the parameters used in the measurements shown in Fig. 4.
@@
We have experimentally investigated the Fraunhofer diffraction pattern of aperiodic slit arrays. In our setup, a fibre-coupled laser beam with $\lambda=660$nm [Thorlabs S1FC660] is first expanded and sent to a spatial light modulator (SLM). The SLM [HoloEye Pluto NIR-015] is programmed to generate the field given in Eq. (1) that is then Fourier-transformed using a lens system with effective focal length f=200f=200mm. The Fraunhofer diffraction pattern is then registered using a Complementary Metal-Oxide-Semiconductor (CMOS) imaging sensor [Basler ace U acA5472-17um]. A sketch of our experimental setup is shown in Fig. 1-(a) (see caption for details).
The chosen slit array has an aperiodic slit distribution described by Eq. (20) and the amplitude distribution of the light going through the slit array is chosen to be the Gaussian function given in Eq. (8) centred around two opposite values of the slit number: $\bar{n}$ and $\bar{n}$. Figs. 1-(b)-(d) present examples of gratings with slit positions following Eq. (20) (on the left), alongside their respective Gaussian intensity distributions (on the right). At first glance, the slit distributions of the illuminated parts may appear periodic, as the aperiodicity becomes more noticeable when analysing a larger portion of the gratings. However, the following experimental results show that the far-field diffraction patterns differ significantly from those of periodic gratings.
@@
Table 1: Far-field periodicities produced by the slit array given in Eq. (20): symbolic expressions presented in Eqs. (21)-(23), values calculated from experimental parameters, and values obtained from the fitting of the measurements shown in Fig. 2.
Symbolic expressions Calculated (mm) Fitted Fig. 2-(a) (mm) Fitted Fig. 2-(b) (mm) Fitted Fig. 2-(c) (mm)
$L_{0}^{\prime}$ $\frac{2\pi}{\xi L}\bar{n}^{-\frac{1}{2}}$ $0369\pm 0.0007 0.0$$0370\pm 0.0007$
$L_{1}^{\prime}$ $\frac{2\pi}{\xi L}2\bar{n}^{\frac{1}{2}}$ $1.48\pm 0.01$ $1.49\pm 0.01$ $1.48\pm 0.01$ $1.47\pm 0.01$
$L_{2}^{\prime}$ $\frac{2\pi}{\xi L}8\bar{n}^{\frac{3}{2}}$ $118\pm 1$
@@
Device resolution, such as the pixel size of modulators and cameras, sets the utmost limit when structuring or measuring light fields. For example, in our experimental setup, the distance between any pair of slits is resolution-limited by the pixel size of our SLM (we use the pixel size as the uncertainty for the length parameter $L$ of the slit array produced). Unlike this experimental limitation, in Subsection 3.1 we have shown that designing an arbitrary aperiodic array with geometry satisfying Eq. (17) for $j=1$ leads to a fundamental resolution limitation. The suppression of the maxima associated with the first-order periodicity by the zeros of the sinc function requires the width of the slits to be greater than the distance between some of the illuminated slits in the array. This condition implies that the part of the array with merged slits loses the aperiodic structure that is responsible for the Fraunhofer diffraction patterns studied in this paper. The resulting aperture is then described by a set of separate slits distributed aperiodically and a large aperture produced by a number of merged slits, as illustrated in Fig. 1-(d).
@@
Figure 3: Experimental results obtained using a grating with slit distribution given by Eq. (20) with $(L=0.400\pm 0.008)$mm and slit width s=0.04s=0.04mm (corresponding to 5 pixels of our SLM). The images correspond to the Fraunhofer diffraction patterns obtained by adjusting the illumination coefficients $A_{n}$ as the Gaussian function, Eq. (8), with $\sigma=\sqrt{2}$ and (a) $\bar{n}=\pm 10$, for (b) $\bar{n}=\pm 17$, and (c) $\bar{n}=\pm 25$. In (d), (e) and (f), we plot the graphs resulting from the sum of recorded images (a), (b) and (c), respectively, over 80 pixels in the vertical direction. Each graph was normalised so that the area under the curve equals one and then divided by the maximum value among all three graphs. The black-dashed line corresponds to the plot of the sinc function described by Eq. (7)
Another way to approach the upper limit of (24) is by increasing the width of the slit $s$ while maintaining fixed values for both $L$ and $\bar{n}$. In this case, $L_{1}^{\prime}$ remains constant, while the positions of the zeros of the sinc function decrease. In Fig. 4 we present the experimental results for a grating with $L=(0.560\pm 0.008)$$m$, $\bar{n}=\pm 25$ and s=0.024s=0.024mm [Fig. 4-(a)], s=0.040s=0.040mm [Fig. 4-(b)] and s=0.056s=0.056mm [Fig. 4-(c)], with the third case corresponding to total suppression of the diffraction maxima at $x^{\prime}=\pm L_{1}^{\prime}$. The respective sums over the vertical direction are shown in Figs. 4-(d), -(e) and -(f). In this case, the reduction of the lateral maxima causes the central peak to become visibly taller. The corresponding aperiodic gratings used in this experiment are illustrated in Figs. 1-(b), (c) and (d).
Figure 4: Experimental results obtained using a grating with slit distribution given by Eq. (20) with $L=(0.560\pm 0.008)$$m$, a Gaussian illumination distribution with $\bar{n}=\pm 25$ and $\sigma=2\sqrt{2}$ The images correspond to the Fraunhofer diffraction patterns obtained with three different slits widths: (a) s=0.024s=0.024mm (3 pixels of the SLM); (b) s=0.040s=0.040mm (5 pixels of the SLM); (c) s=0.056s=0.056mm (7 pixels of the SLM). In (d), (e) and (f), we plot the graphs resulting from the sum of recorded images (a), (b) and (c), respectively, over 80 pixels in the vertical direction. Each graph was normalised so that the area under the curve equals one and then divided by the maximum value among all three graphs. The dashed black line corresponds to the plot of the sinc function described by Eq. (7). The corresponding aperiodic gratings used in this experiment are illustrated in Figs. 1-(b), (c) and (d).
@@ 5 Conclusions
We have theoretically and experimentally investigated the Fraunhofer diffraction patterns produced by aperiodic slit arrays. We provide general conditions for the observation of quasiperiodic oscillations over different length scales based on the geometry and illumination pattern of the array. Most interestingly, aperiodic diffracting structures with monotonically decreasing distance between adjacent slits suffer from a fundamental resolution issue: there comes a point in which the separation of the slits must be smaller than the width of the slits themselves. In this work, we have shown that adjustment of the grating geometry to suppress the peaks of the diffraction maxima associated with its aperiodicity necessarily requires some of the slits to merge into a completely opened aperture. These results provide new insights into the far-field diffraction properties of aperiodic gratings, with possible applications in beam shaping, the design of diffractive devices, and the characterisation of materials with quasiperiodic structures. Finally, we note that in this work we have not discussed the near-field diffraction patterns generated by these aperiodic structures, which feature distorted Talbot patterns that may be of independent interest (for an example, see Ref. [20], Fig. 3-(c)). We leave this discussion for future work.
@@
The data supporting this study are available in the Zenodo repository at DOI: 10.5281/zenodo.18476750.